### v1

@@ Introduction
For a brief overview of the current state of the Odderon problem, see, for example, [1].
In our note we are more specific and would like to touch upon one extreme case of realization of the Odderon as a ”flat Regge trajectory” [2] (actually a fixed pole) with help of which several ”two-channel” eikonal model variants were tested with a data description on $pp$-scattering at $\sqrt{s}=50\div 13000$ GeV and for $-t<0.1\;GeV^{2}$. With use of some special prescription for calculation of the $\chi^{2}$ the authors of the paper in question claim quite a reasonable description of the data with $\chi^{2}/ndf=1.11$ and indicate that the absence of the Odderon contribution worsens the quality of the description though this contribution is small. The sign of the Odderon was also defined and the value of the Odderon coupling $\beta_{\mathcal{O}}(t)$ at $t=0$ is estimated to be less than that of the Pomeron. Moreover, the contribution of the Odderon to the ratio $\Re T^{pp}(s,0)/\Im T^{pp}(s,0)$ was found to be an order of magnitude smaller than that proclaimed by the TOTEM collaboration.
@@
$=\pi\alpha^{{}^{\prime}}(t)\left[\alpha(t)+\frac{1}{2}\right]\left\{i+\mbox{tg}\left[\frac{\pi\alpha(t)}{2}\right]\right\}\Gamma_{\alpha(t)}^{2}(t)P_{\alpha(t)}(-z_{t})$
@@
The propagator stems from $tg(\pi\alpha(t)/2).$ There are no poles (physical states) in the cross-channel of amplitudes of the ”alien” (in this case positive, with even $J$) signature.
@@
at low $t$ and then monitor the consequences. Taking afterwards $\alpha^{{}^{\prime}}=0$ we reproduce the ”flat Odderon” premise of Ref.[2]. Substiting $\alpha(t)\rightarrow 1+\alpha^{{}^{\prime}}t$ in Eq.(1) and considering it in the limit $\alpha^{{}^{\prime}}\rightarrow 0,t\rightarrow 0$ we get
@@
So we see that at the level of the Born amplitude the Odderon decouples from the proton at $t=0$. The real part decouples no weaker than tsimilar-$\sim t$, while the imaginary part decouples no weaker than $\sim t^{2}$.
@@
It is important to note that authors provide both the central values of the fitted parameters of the models and their errors. However, the parameters extracted from the models (for example, the parameter $\rho$ [6] ) are provided only with their central values; for some reason, the authors do not provide the errors which can be of a paramount importance for the final conclusions. Namely, in the experimental data at the $t$ intervals the values of $\mid t\mid$ were carefully selected based on the high confidence level (at least $80\%$) for the extracted parameters. In particular, the work [5] shows that $\rho=0.01\pm 0.04$, i.e. the extracted value is determined with an accuracy of $40\%$(p. 054003-9). This is fundamentally at odds with the error of the authors of the TOTEM experiment, who groundlessly claim an error in the $\rho$parameter to be only $10\%$(with the same central value). It is clear that with such an error value, it is at least premature to talk about a discrepancy with the predictions of the COMPETE collaboration[7]. So, arguing about the difference in the values of the parameters in this work is generally pointless without indication of the errors. The authors of [2] refer to the work just mentioned [5] , but they cite intermediate values of the parameter $\rho$ from this work, and not its final value and its error from the conclusions of the referenced article. Moreover, the authors, referring to work[5], fairly note that its results were obtained on the basis of the TOTEM experiment only. Just because the official results of the ATLAS experiment did not yet exist at that time.
@@
This allows us to calculate the confidence level of this extraction of parameters. So, in more than half of the cases, this confidence level is extremely low or significantly lower than $50\%$. On such a background claims that ”the presence of the C-odd(Odderon)contribution essentially improves” the quality of the data description seem speculative and unfounded.

### v2

@@ Introduction
For a brief overview of the current state of the Odderon problem, see, for example, [1], [2] .
In our note we are more specific and would like to touch upon one extreme case of realization of the Odderon as a ”flat Regge trajectory” [3] (actually a fixed pole) with help of which several ”two-channel” eikonal model variants were tested with a data description on $pp$-scattering at $\sqrt{s}=50\div 13000$ GeV and for $t<0.1\;\mbox{GeV}^{2}$. With use of some special prescription for calculation of the $\chi^{2}$ the authors of the paper in question claim quite a reasonable description of the data with $\chi^{2}/ndf=1.11$ and indicate that the absence of the Odderon contribution worsens the quality of the description though this contribution is small. The sign of the Odderon was also defined and the value of the Odderon coupling $\beta_{\mathcal{O}}(t)$ at $t=0$ is estimated to be less than that of the Pomeron. Moreover, the contribution of the Odderon to the ratio $\Re T^{pp}(s,0)/\Im T^{pp}(s,0)$ was found to be an order of magnitude smaller than that proclaimed by the TOTEM collaboration.
@@
$\pi\alpha^{{}^{\prime}}(t)\left[\alpha(t)+\frac{1}{2}\right]\left\{i+\mbox{tg}\left[\frac{\pi\alpha(t)}{2}\right]\right\}\Gamma_{\alpha(t)}^{2}(t)P_{\alpha(t)}(-z_{t})$
@@
The propagator stems from $\mbox{tg}[\pi\alpha(t)/2].$ There are no poles (physical states) in the cross-channel of amplitudes of the ”alien” (in this case positive, with even $J$) signature.
@@
at low $t$ and then monitor the consequences. Taking afterwards $\alpha{{}^{\prime}}=0$ we reproduce the ”flat Odderon” premise of Ref. [3]. Substiting $\alpha(t)\rightarrow 1+\alpha{{}^{\prime}}t$ in Eq.(1) and considering it in the limit $\alpha{{}^{\prime}}\rightarrow 0$, $t\rightarrow 0$ we get
@@
So we see that at the level of the Born amplitude the Odderon decouples from the proton at $t=0$. The real part decouples no weaker than $\sim t$, while the imaginary part decouples no weaker than $\sim{t}^{2}$.
@@
However, first, let us make one important remark. There is extensive literature on the analysis of diffraction scattering parameters at various energies. However, in this paper, we rely primarily on experimental data at 13 TeV for the following reason: in our view, the measurements at 13 TeV are unique in that they were first performed on the same beam and at the same energy using two different instruments – ATLAS and TOTEM. These data turned out to differ significantly. As a consequence of this fact, the extracted theoretical parameters differ significantly, and the authors of numerous papers point to a variety of reasons for their discrepancies. For example, in [12], the authors point to an underestimation of systematic errors at the TOTEM setup and the possibility of a slower increase in the total cross section in the pp interaction. A significant caveat is noted in the Pierre Auger Observatory report [10] which emphasizes the preference for the ATLAS total cross-section data over the TOTEM data.
There are numerous such papers examining the possible causes of these different experimental data (see, e.g. [11]). However, the authors of the ATLAS and TOTEM experimental setups remain silent on the reasons for such divergent measurements of the same differen tial cross-section at 13 TeV. A joint analysis of the sources of these different results, which are a result of the design of these experimental setups, would be fundamentally important. As a consequence of this state of affairs, we are forced to rely solely on published experimental data from these setups, which significantly contradict each other. And what to do with these contradictions in the absence of an analysis by their creators is anyone’s guess.
There is no doubt that these measurements provide approximately correct values for the parameters of various theoretical models, but they are fundamentally unsuitable for sifting through different theoretical concepts (calibrating theories). For this reason, we rely only on published experimental results from the ATLAS and TOTEM facilities at 13 TeV and do not delve into the details of their differences.
The obvious, significant differences in experimental measurements at 13 TeV, identified for the first time, also give reason to doubt the measurements at other energies. In essence, this assertion is the main content of our note. For this reason, we do not analyze the sources of the various measurement results, of which there are already a large number.
It is important to note that authors provide both the central values of the fitted parameters of the models and their errors. However, the parameters extracted from the models (for example, the parameter $\rho$ [7] ) are provided only with their central values; for some reason, the authors do not provide the errors which can be of a paramount importance for the final conclusions. Namely, in the experimental data at the $t$ intervals the values of $|t|$ were carefully selected based on the high confidence level (at least $80\%$) for the extracted parameters. In particular, the work [6] shows that $\rho=0.1\pm 0.04$, i.e. the extracted value is determined with an accuracy of $40\%$ (p. 054003-9). This is fundamentally at odds with the error of the authors of the TOTEM experiment, who groundlessly claim an error in the $\rho{-}$parameter to be only $10\%$ (with the same central value). It is clear that with such an error value, it is at least premature to talk about a discrepancy with the predictions of the COMPETE collaboration[8]. So, arguing about the difference in the values of the parameters in this work is generally pointless without indication of the errors. The authors of [3] refer to the work just mentioned [6], but they cite intermediate values of the parameter $\rho$ from this work, and not its final value and its error from the conclusions of the referenced article. Moreover, the authors, referring to work[6], fairly note that its results were obtained on the basis of the TOTEM experiment only. Just because the official results of the ATLAS experiment did not yet exist at that time.
@@
Table III
Ensemle A
DoF (nu\ nu) 332 332 332 332 332
$\chi^{2}/\nu$ 0.96 0.97 0.97 0.97 0.96
CL [%] 69.2 64.4 64.4 64.4 69.2
Ensemle T
DoF ($\nu$) 418 418 418 418 418
$\chi^{2}/\nu$ 1.28 1.30 1.29 1.28 1.27
CL $\cdot 10^{3}$ 6.8 2.3 4.0 6.8 11.3
Ensemle A $\oplus$ T
DoF ($\nu$) 504 504 504 504 504
$\chi^{2}/\nu$ 1.11 1.12 1.11 1.10 1.09
CL [%] 4.3 3.1 4.3 5.9 7.9
Figure 1: The values of the confidence levels in Table III (see [3]).
@@
These results from the authors’ fittings [3] allow us to calculate the confidence level of their parameter extraction results. We have supplemented Table III of the authors’ table with rows detailing the confidence levels of these results – see Fig. 1.
The supplemented table shows that supplementing the general data set with ATLAS data at 13 TeV yields a decent confidence level—over 50(Ensemple A). Meanwhile, a similar supplement with TOTEM data at 13 TeV reduces the confidence level to practically zero (Ensemple T). Clearly, drawing conclusions after this that the theoretical parameters extracted from one array are preferable to the same parameters extracted from another array is pointless. Against this background, claims that ”the presence of a C-odd (odderon) contribution significantly improves” the quality of the data description appear speculative and unfounded. And even more so, combining the overall data set by simultaneously adding data from the ATLAS and TOTEM experiments at 13 TeV (Ensemle A $\oplus$ T) is also an exercise devoid of any meaning.